Investigators reported the resuls of a study designed to assess whether or not St. Johns Nort is effective in treating m cases of depression. The study involved 338 subjects, randomly assigned to receive one of three treatments. TH results are presented below: St. Johns Wort Placebo Zoloft Total Full Response 27 Partial response 16 No response 70 Total 37 91 27 26 56 192 109 338 13 113 116 What is the probability that a randomly selected subject had no response or was treated with Zoloft? O a Ans. not listed b. 791338 с. 271338 O d. 43/113 e. 116/338 O f. 79/247 O9-66/338 h. 245/338 . 27/109
Answer #1

1)P(no response or was treated with Zoloft) =(192+109-56)/338 =245/338

option H is correct